Find the area of the region inside both the rose curve r=sin⁡〖(2θ)〗 and the circle r=cos⁡θ. Choose the most appropriate graph and intervals to the best of your interest to solve this question
Answer #1

1) Given 8=sino, &- coso 8=sin2o O=T/6 8=coso sin2o=coso 2sino coso -COSO =D COSO (25ine -1 - 0 Coso =0, asino-1=0 Sino = % o